OpenStudy (anonymous):

Write an equation of a line that passes through the point (2, 4) and whose slope is ½.

OpenStudy (anonymous):

The eq of line in slope point form is give as \[y-y_{1}=m(x-x_{1}) \rightarrow y-4=\frac{1}{2}(x-x_{1}) \] Multiplying both sides by 2 we find \[2y-8=x-2 \rightarrow 2y= x+6 \rightarrow -x+2y=6 \rightarrow x-2y=-6 \] Which is the required eq of line.